# Tree-tubings and the combinatorics of resurgent Dyson–Schwinger equations

### v1

@@ 4.2. Transseries approach
where higher order contributions in $\sigma_{1},\ldots,\sigma_{m}$ can be computed mechanically (see, e.g., [5, 6, 4]).
@@
For $m=1$ the formal solution begins as
@@
For $m=2$, with $a_{1}<a_{2}$, the formal solution begins as
@@
Experimentally, we find that the procedure described an the beginning of Section 4.2 still works if we modify our transseries ansatz to have the following shape: [^4] The precise shape of this ansatz can be deduced from a Frobenius analysis of the second order linear homogeneous ODE satisfied by $\gamma_{\rm trans}$.
@@
Experimentally, working with some thousands of terms generated with the gamma.py program (see Remark 3.7), it appears that the large-order asymptotics of the perturbative expansion coefficients changes abruptly when $a_{1}=a_{2}$; another reflection of the fundamental difference between the behaviour of the solution when $a_{1}=a_{2}$ vs $a_{1}\neq a_{2}$.
We can give a combinatorial heuristic for why the $a_{1}=a_{2}$ case ought to have different growth behavior. However, note that this could only be part of a rigorous argument in the special case of the leaf tubings, see Subsection 3.1.3. The heuristic is as follows: any difference in behavior depending on properties of the values of the $a_{i}$ should be seen combinatorially in the compatible labellings. The effect of the compatible labelling constraint is most pronounced when many tubes share a root since that is when the admissibility constraint on words comes into play most strongly. Consequently, for a heuristic as to why $a_{1}=a_{2}$ should behave differently from $a_{1}\neq a_{2}$ we can look to why the number of admissible words behaves differently when $a_{1}=a_{2}$ compared to $a_{1}\neq a_{2}$. For the leaf tubings, all tubes of size $>1$ share a root, and so counting the number of compatible labellings of a leaf tubing of a tree on $n+1$ vertices is exactly counting the number of admissible words of length $n$. For other tubings, we do not have an exact count for the number of compatible labellings, but we can still use the number of admissible words as a useful heuristic for what kinds of differences we might see.
@@
Example 4.8 (Third order case).
For $m=3$, with $a_{1}<a_{2}<a_{3}$, the formal solution begins as
@@
For the $\phi^{3}$ example, we have $\{a_{1},a_{2},a_{3}\}=\{2,3,6\}$, and the formal solution is: $\gamma_{\rm formal}(x)=x+11x^{2}+376x^{3}+20241x^{4}+\dots$, [17, A051862]. The 3 independent non-perturbative contributions are of the shape [6]

### v3

@@ 4.2. Transseries approach
where higher-order contributions in $\sigma_{1},\ldots,\sigma_{m}$ can be computed mechanically (see, e.g., [8, 9, 7]).
@@
For $m=1$ the formal solution begins as follows:
@@
For $m=2$, with $a_{1}<a_{2}$, the formal solution begins as follows:
@@
Experimentally, we find that the procedure described at the beginning of Section 4.2 still works if we modify our transseries ansatz to have the following shape: [^4] The precise shape of this ansatz can be deduced from a Frobenius analysis of the second-order linear homogeneous ODE satisfied by $\gamma_{\rm trans}$.
@@
We do not know if such a degenerate case, where the values of two $a_{i}$ coincide and $F(\rho)$ has a pole of order two, is allowed by the various constraints that quantum field theory imposes on $F(\rho)$. However, we can provide an interpretation of this degeneracy and the different asymptotics it produces on the combinatorial side:
@@
Experimentally, working with some thousands of terms generated with the gamma.py program (see Remark 3.8), it appears that the large-order asymptotics of the perturbative expansion coefficients changes abruptly when $a_{1}=a_{2}$; another reflection of the fundamental difference between the behavior of the solution when $a_{1}=a_{2}$ vs $a_{1}\neq a_{2}$.
We can give a combinatorial heuristic for why the $a_{1}=a_{2}$ case ought to have different growth behavior. However, note that this could only be part of a rigorous argument in the special case of the leaf tubings, see Subsection 3.1.3. The heuristic is as follows: any difference in the behavior depending on properties of the values of the $a_{i}$ should be seen combinatorially in the compatible labelings. The effect of the compatible labeling constraint is most pronounced when many tubes share a root since that is when the admissibility constraint on words comes into play most strongly. Consequently, for a heuristic as to why $a_{1}=a_{2}$ should behave differently from $a_{1}\neq a_{2}$ we can look to why the number of admissible words behaves differently when $a_{1}=a_{2}$ compared to $a_{1}\neq a_{2}$. For the leaf tubings, all tubes of size >1>1 share a root, and so counting the number of compatible labelings of a leaf tubing of a tree on $n+1$ vertices is exactly counting the number of admissible words of length $n$. For other tubings, we do not have an exact count for the number of compatible labelings, but we can still use the number of admissible words as a useful heuristic for what kinds of differences we might see.
@@
Example 4.8 (Third-order case).
For $m=3$, with $a_{1}<a_{2}<a_{3}$, the formal solution begins as follows:
@@
For the $\phi^{3}$ example, we have $\{a_{1},a_{2},a_{3}\}=\{2,3,6\}$, and the formal solution is: $\gamma_{\rm formal}(x)=x+11x^{2}+376x^{3}+20241x^{4}+\dots$, [21, A051862]. The 3 independent non-perturbative contributions are of the shape [9]
@@
An interesting feature of this $\phi^{3}$ model is the appearance of logarithmic terms [9, 7] at higher orders in the transseries for the anomalous dimension $\gamma(x)$, due to the resonance effect: $\{\frac{1}{a_{1}},\frac{1}{a_{2}},\frac{1}{a_{3}}\}=\{\frac{1}{2},\frac{1}{3},\frac{1}{6}\}=\{3,2,1\}\times\frac{1}{6}$. We leave to future work the interesting combinatorial question of investigating the general conditions for such logarithmic terms to be generated.